Assignment > Open Assignment PRINTER VERSION BACK NEXT Chapter 5, Section 5.3, Question 03 Determine 4 (xo), (x0) and 6Assignment > Open Assignment B PRINTER VERSION BACK NEXT Equation Editor Common S Matrix sinia) cola) tana) ASSIGNMENT RESOUR
Answer #1

þ(No) +622 6 (Ho) + sin no) $ (20)=0 , 0(X) + 6 to y( Mo) + sin(no). Y CO )=0 As, Ho=o im this IVP: $ (Mo) + 6. 07. y(oger . ( Mo) +62 (M.) + (12.10 + sin(no)) $(10) + (ros no ) $(Mo) = 0 i ! (no) t boot. f(Ho) + (12.00 + sino y/10) + (cos(